From: STOSC::SMM "On-Duty SMM x5074" 7-JAN-1993 13:40:34.30 To: @[.DIS]HST_STAT CC: Subj: HST Status Report, 01/06 - 01/07. HUBBLE SPACE TELESCOPE DAILY REPORT #798 PERIOD COVERED: 0800 EST 01/06/93 - 0800 EST 01/07/93 Daily Status Report as of 007/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/BL 4057 (Spectropolarimetry of QSOs) The Faint Object Spectrograph (Blue Side) was used to make three spectropolarimetry observations of Markarian galaxy MARK421 below 3000 A0. All three observations were received and displayed in the OSS. They were reported to look okay. 1.2 Completed WFC 4105 (HST Medium Deep Survey, High Latitude) The Wide Field Camera was used to take two high latitude image in parallel with the prime science observations. Both image were received and displayed. They both were reported to look okay. 1.3 Completed FOS/BL 4495 (GIMP Test) The Geomagnetic Induced Motion Problem (GIMP) test was conducted on the correction mechanism for the Faint Object Spectrograph. All observations were received and displayed in the OSS. The four rapid mode wave length calibration, and the 24 image mode TALEDs were all reported to be okay. 1.4 Completed WFC-PC 4109 (Rapid Internal Monitor) Both the Wide Field Camera and the Planetary Camera were used to take two internal flat images looking for new forms of contamination (i.e., daisies and measles). Both images were received and displayed. the images were reported to be okay. 1.5 Completed Four Sets of PC 3172 (Internal Monitor Proposal) The Planetary Camera was used to take a series of bias, dark, internal flat, and KSPOT images. No external target was used. The dark observation looked fine. Of the Internal flat observations: one had some saturation in the corners, one was underexposed, and one was reported to be okay. The KSPOT observation was reported to have some internal reflections from the grossly overexposed Baum spot. All other observations were reported to be okay. 1.6 Completed PC 4084 (Globular Cluster Cores, GTO Augmentation) The Planetary Camera was used to make two images of the core of globular cluster BD+75D325. Both images were executed in course track. The point spread function of the test star was well placed in CCD PC #6. 1.7 Completed PC 4347 (Flood Retention Test) The Planetary Camera was used to obtain three Earth flat calibration images using filter F439W. All three images were received and displayed. The first two images were reported to be saturated. Part of the third observation was lost due to a track change. 1.8 Completed HSP/UV2/POL 3234 ( X-Ray Binaries) The High Speed Photometer was used to obtain both photometric and polarimetric light curves of the x-ray binary star HD77581 (aka, Vela XR-1). The observations were executed in fine lock. Both image were reported to be well centered and to look okay. 1.9 Completed WFC 1306 (GTO Low Latitude Parallel Program) The Wide Field Camera was used to take low latitude images in parallel with the prime science observations. The observations were executed as scheduled, and no problems were reported. 1.10 Completed WFC 3605 (Stellar Winds and Post-Main Sequence Evolution) The Wide Field Camera was used to make observations of star BE294 (Spectral Class Ofpe/WNL) which has a visual magnitude of 11.9, and is located in the Large Magellanic Cloud. This star is known to be very hot and luminous, and have a mass loss rate much higher than normal O class stars. The observations were executed as scheduled, and no problems were reported. 1.11 Initiated HRS/2 3825 (Novae in Outburst in the Magellanic Clouds) The High Resolution Spectrograph was used to make target of opportunity spectral observations of novae in the Large Magellanic Cloud during their nebular phase. These observations were initiated as scheduled, and are currently under way. No problems have been reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 3 (007/0554 and 0715Z) HSTAR written Scheduled Reacquisitions : 7 Successful : 7 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 8 Successful : 8 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete GHRS/@ 3825 (Novae in Outburst in the Magellanic Clouds) 3.2 complete FOS/BL 4495 (GIMP Test) 3.3 Uplink Updated Ephemris Tables, Per MOSES 2871 3.4 Complete Two Sets of FOC/96 3801 (Search for QSOs suitable for Observation) 3.5 Complete HSP/UV2/POL 3234 (X-Ray Binaries) 3.6 Complete FOS/BL 4212 (Internal Wave Length Calibrations) 3.7 Initiate GHRS/2 4217 (Absorption Cross Sections of Nearby Galaxies) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time -END-